# Spin Theory Based on the Extended Least Action Principle and Information Metrics: Quantization, Entanglement, and Bell Test With Time Delay

### v2

@@ V.3 Other Bell States
On the other hand, if the correlations between the electron pair are such that $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, and $\theta_{A}^{y}=\theta^{y}_{B}$ along the $y$-axis, we must also have
@@
Thus, measurement along the $y$-axis results in the same spin directions for both electrons, while measurement along the $z$-axis yields opposite spin direction. A pair of electrons described by both (63) and (66) shares the same correlation specified by the spin triplet states in standard quantum mechanics
@@
Similarly, we can give the equivalent descriptions of the other two Bell states in standard quantum mechanics. In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, one has $\varrho(\Theta^{+}_{A})=\varrho(\Theta^{-}_{A})=1/2$, (60) becomes
$\tilde{p}(\theta_{A}^{z},\theta_{B}^{z})=\frac{1}{2}\delta(\theta_{A}^{z})\delta(\theta_{B}^{z})+\frac{1}{2}\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z}-\pi)$ (68)
@@ V.4 Bell Inequality
Bell inequality is developed to prove that local hidden variable theory cannot produce the predictions of standard quantum mechanics. Bell inequality is violated when tested with an entangled pair of photons or electrons, demonstrating there is non-local correlation in standard quantum mechanics. Even though modern Bell experiments Hensen confirm that such non-local correlation does not imply non-local causality, such Bell non-locality is still daunting to the quantum physics community because there is no intuitive physical picture to explain the phenomenon. Here we will investigate what insights the present theory can bring to this challenge.
@@
Here $a,a^{\prime}$ are detector settings for Alice, and $b,b^{\prime}$ are detector settings for Bob. They usually refer to the directions of the magnetic field in the spin measurement and can be considered as unit vectors for the corresponding directions. The term $E(a,b)$ is the expectation value of measurement outcomes. That is, the statistical average of $S_{A}(a)S_{B}(b)$ where $S_{A}(a)$ and $S_{B}(b)$ are the spin measurement outcomes for setting $a$ and $b$, respectively. For the setting $a$, $S_{A}(a)=+1$ for spin-up and $S_{A}(b)=-1$ for spin-down. Similar meanings can be inferred for the other $E(\cdot)$ terms in (71).
The expectation value $E(\cdot)$ depends on the state of the entangled pair. Suppose the entangled electron pair is in the singlet described by the joint probability density (63) and (64). It is important to note that both (63) and (64) are needed to give a complete description of the singlet state. In Appendix D, we prove that given the joint probability density (63) and (64),
@@
The reason for the violation of the CHSH inequality is due to the fact that the joint probability density (63) and (64) cannot be factorized. In other words, they cannot be written as a product of two factors, one only contains variable for $A$ and the other only contains variable for $B$. Such correlation can be maintained even though the two electrons are remotely separated. The root cause of such correlation is that the random variables are correlated, $\theta^{z}_{A}+\theta^{z}_{B}=\pi$ and $\theta^{y}_{A}+\theta^{y}_{B}=\pi$. These random variables are intrinsically local to the electrons. The term non-local correlation, or Bell non-locality, is misleading. Instead, as Hall pointed out Hall2015 , a better term is Bell non-separability.
Figure 7: Bell test with time delay. Starting with a Bell test experiment setup that has already successfully confirmed the violation of Bell-CHSH inequality. Then, we add a time delay in the measurement. Bob delays his measurement by a time period $\Delta t$, while the signals from Alice’s Stern-Gerlach detectors are delayed by the same $\Delta t$ before it is sent to the coincident monitor. When $\Delta t$ is chosen to be sufficiently large, we expect the Bell-CHSH inequality to become non-violated.
@@ VI Possible Experiments to Confirm the Spin Theory
The spin theory presented here so far recovers many of the results from standard quantum mechanics. However, when these conditions are not taken to the limits, the present theory will give some results different from standard quantum mechanics. In this section, we will provide two possible experiments to confirm the difference.
@@ VI.1 Dependency of Quantization on Interaction with Magnetic Field
One of the crucial assumption in the present theory is the derivation of (15), which assumes that the parameter $m$, which is related to the order of Tsallis relative entropy as shown in (12), monotonically increases when the electron travels along an inhomogeneous magnetic field $B_{z}$. Standard quantum mechanics postulates that quantization of electron spin is an intrinsic property. Practically, it corresponds to the case that whenever there is an external field $B_{z}$ is applied, quantization occurs instantaneously, such that only two discrete results can be observed. However, the theory presented here allows $m$ taking a finite number when the magnetic field $B_{z}$ is applied. In that case, the probability density (13) should be used instead of the probability density (15) with two quantized outcomes.
@@
where $^{B}_{0}\hat{z}$ is a constant field, and the coefficient $\eta$ determines the strength of the magnetic field. Since the orientation of the intrinsic magnetic moment of the electron follows a probability density distribution (13), the electron passing the Stern-Gerlach apparatus will experience a force that also depends on the orientation. Between angles $\theta$ and $\theta+d\theta$, the force is
@@
The above calculation ignores the effect of the translational component of vacuum fluctuations and calculates the displacement along the $z$-axis using classical mechanics. In theory, this approximation is reasonable when $\Delta T$ is sufficiently large so that $\Delta z$ is much larger than the effect of the translational component of vacuum fluctuations. Eq. (75) shows that in a magnetic field with extremely small $\eta$, the electrons passing the Stern-Gerlach apparatus will reach the detector screen with a continuous distribution, as shown in Fig. 2b.
However, as $\Delta T$ increases, parameter $m$ monotonically increases as well, the measurement outcomes may rapidly converge to two discrete lines, and yield the same prediction as standard quantum mechanics. The possibility to set up an experiment to test the deviation from standard quantum mechanics depends on the balance of two factors, 1.) choosing the small $\eta$, and 2.) choosing a proper time duration $\Delta T$ that determines how long the electron is interacting with $\vec{B}$. Such an experiment can be challenged to implement. The experiment proposed next can be more realistic.
@@ VI.2 Bell Test with Time Delay
In Section V.1, we show that for an entangled electron pair described by the joint probability density (68) and (64), the entangled relation can be maintained without time constraint. However, when Alice performs her measurement on the spin of electron $A$, obtains spin-up, and makes her prediction on Bob’s measurement result of spin-up along the same direction, the prediction is only valid for a limited time. That is, there is a time constraint specified in (61) that Bob needs to perform his measurement on the spin of electron $B$ in order to confirm Alice’s prediction. If $t_{m}^{B}>t_{m}^{A}+\Delta t^{B}_{+}$, Bob will obtain the measurement result of spin-down instead. Furthermore, a singlet spin state is described by two probability density functions (63) and (64). After Alice performs her measurement on electron $A$ and obtains result of spin-up, if Bob delays his measurement by a time such that the correlation described by either (63) or (64) is no longer satisfied, the singlet state is degraded and no longer valid.
@@
We can explore the effect of entanglement degradation for a singlet state due to Bob’s delayed measurement in the Bell test experiment, as illustrated in Figure 7. In a Bell test experiment, the electron source generates $N$ copies of entangled electron pairs in a singlet state. For each pair, electrons $A$ and $B$ are sent to two Stern-Gerlach apparatuses that are remotely separated [^7] As mentioned earlier, in modern Bell test experiment Hensen , the two electrons are already separated remotely before being prepared to be entangled. The electron entanglement is achieved by entanglement swapping with photons. There is no need to move the electrons and therefore, no need to worry about the impacts on the correlation due to the movement of electrons.. Alice measures the spin of electron $A$ along an axis labeled by unit vector $a$. At the same time, Bob measures the spin of electron $B$ along an axis labeled by unit vector $b$. The detectors for each Stern-Gerlach apparatus detect the spin-up and spin-down results, and generate corresponding signals that are sent to the coincidence monitor. Counting the four types (++,+,+,)(++,+-,-+,--) of coincidences for the $N$ copies of the singlet spin pairs, one can calculate whether the Bell-CHSH inequality is violated. As shown in Section V.4, for the singlet spin pair described by (63) or (64), if Alice chooses $a=0$ and $a^{\prime}=\pi/2$, and Bob chooses $b=\pi/4$ and $b^{\prime}=3\pi/4$, maximum violation of the Bell-CHSH inequality is achieved.
@@
Using exactly the same experiment setup, we add an extra step here. Instead that Bob performs the measurement on electron $B$ at the same time as Alice’s measurement, he delays his measurement by a time period of $\Delta t$. When $\Delta t$ is sufficiently large, the entanglement correlation of a singlet state is degraded, and we expect that the Bell-CHSH inequality will not be violated. Since we do not know how long to delay in order to observe the non-violation of Bell-CHSH inequality, we can start with a very small delay $\Delta t_{1}$ such that the Bell-CHSH inequality is still violated, and then gradually increase the delay. We repeat the experiments $M$ times, but choose the delay times such that $\Delta t_{1}<\Delta t_{2}<\ldots<\Delta t_{M}$. Each experiment consumes $N$ copies of singlet electron pairs. According to the present theory, we expect the violation of Bell-CHSH inequality will be gradually reduced, and at some time the results will satisfy the Bell-CHSH inequality. The reason for this is that with a sufficient large delay $\Delta t$, the orientation of intrinsic angular momentum of electron $B$ can fluctuate to switch from trending upward to trending downward, or vice versus, as depicted in Figure 4. When a orientation switching occurs, either $\theta_{B}\in\Theta_{B}^{+}\to\theta_{B}\in\Theta_{B}^{-}$, or $\theta_{B}\in\Theta_{B}^{-}\to\theta_{B}\in\Theta_{B}^{+}$, the initial correlation between $\theta_{A}$ and $\theta_{B}$ is changed. As $\Delta t$ increases, the orientation of electron $B$ appears as a random variable again for the $N$ copies of electron pairs, so that the initial correlation between $\theta_{A}$ and $\theta_{B}$ is no longer hold. Consequently, the Bell-CHSH inequality becomes satisfied. Appendix D gives the mathematical calculation of such a scenario.
@@
The reason the entanglement relation between $A$ and $B$ is degraded is due to the orientation fluctuations of the intrinsic angular momentum. This is different from the decoherence theory where the entanglement relation between $A$ and $B$ can be destroyed by environmental disturbance. Here, however, the degradation of entanglement during measurement is intrinsic to the electron pair, nothing to do with the environment. Without the measurement event, the entanglement relation between $A$ and $B$ can be maintained without a time limit. Certainly, if there is also environmental disturbance, the entanglement relation will be destroyed without measurement. So the present theory is not incompatible with the decoherence theory. However, if confirmed by experiment, the intrinsic degradation of entanglement during measurement can have a non-negligible implication on its applications in quantum computing and quantum information.
@@
The challenges to perform the above experiment come from choosing the proper time delay $\Delta t$. To meet the locality condition of Bell test, $\Delta t$ needs to be smaller than the time for a light signal traveling from Alice to Bob. This is to eliminate the potential loophole of hidden signals sent from Alice to Bob. But $\Delta t$ can not be too small either so that the orientation of intrinsic angular momentum of $B$ switches from $\theta_{B}\in\Theta_{B}^{+}\to\theta_{B}\in\Theta_{B}^{-}$, or $\theta_{B}\in\Theta_{B}^{-}\to\theta_{B}\in\Theta_{B}^{+}$. To resolve this contention, one can separate electrons $A$ and $B$ with a very long distance so that the time window to comply with locality condition is large enough. A more subtle argument is that since the goal here is to show that the Bell-CHSH inequality is not violated, the requirement of locality is not applicable. Thus, $\Delta t$ does not need to be smaller than the time for a light signal traveling from Alice to Bob.
@@ VII.1 Conceptual Implications
One advantage of the present theory is it provides an intuitive physical picture to understand the phenomenon of entanglement between two electron spins. As shown in Section V, entanglement is simply the manifestation of the correlation between the orientations of intrinsic angular momenta of two electrons. Even though the orientations are random variables, they can have correlation due to previous interactions between them, or due to previous interactions with common external fields. More importantly, the orientation of intrinsic angular momentum is an intrinsically local property. Once the correlation of orientations between two electrons is established, it can be maintained even if the two electrons are remotely separated, until there is measurement performed on them. Performing measurement on one electron will not cause any change on the other electron. However, the measurement outcome reveals the orientation of the measured electron’s intrinsic angular momentum, which enables us to infer the spin measurement outcome of the other electron because of the correlation on the orientations. The seemingly non-local correlation has nothing to do with any non-local causal relation.
@@
To further clarify the point that there is no non-causal effect in the Bell experiment with a pair of electrons in the singlet state, suppose after Alice performs her experiment and obtains the result of spin-up for electron $A$, she does not send the measurement outcome to Bob. Since the measurement action of Alice is local and has no influence on electron $B$, from Bob’s point of view, he still perceives both $A$ and $B$ in the original singlet state described by (63) and (64). Alice’s and Bob’s descriptions on $B$ are different but both are valid. This is consistent with the relational quantum mechanics (RQM) interpretation advocated by Rovelli Rovelli:1995fv ; Rovelli07 . Since Bob does not know the measurement outcome from Alice, he predicts that if he measures electron $B$ along any direction, he will obtain spin-up and spin-down with equal probability. However, if Alice sends the measurement results to Bob, Bob can infer the orientation of intrinsic angular momentum of electron $B$. Bob’s knowledge on electron $B$ is updated with the new information from Alice. Thus, he can predict if he measures electron $B$ along the same direction as Alice’s measurement, he will obtain the result of spin-down. This analysis also illustrates that the wave function in quantum mechanics is just a mathematical tool and not associated with certain ontic properties.
@@
We see the spin model and theory presented here gives an intuitive physical picture that may help to consolidate several quantum physics interpretations. In addition, our theory supports the idea that measurement outcome is context dependent. Measurement outcome for the spin of an electron depends on the measurement context, that is, the direction of magnetic field used in the Stern-Gerlach apparatus. The two level discreteness of measurement outcome is the result of a sufficiently strong gradient of magnetic field. For an electron with initially completely random orientation of intrinsic angular momentum, we can measure the spin along any direction but always get spin-up and spin-down with equal probability. There is no predefined spin property without specifying an experimental context.

### v3

@@ V.3 Other Bell States
In previous sections, we show that the spin singlet state can be described by (61) and (62). Here, we would like to recover other Bell states using the probability density functions. If the correlations between the electron pair are such that $\theta_{A}^{z}+\theta^{z}_{B}=\pi$ along the $z$-axis, and $\theta_{A}^{y}=\theta^{y}_{B}$ along the $y$-axis, we must also have
@@
Thus, measurement along the $y$-axis results in the same spin directions for both electrons, whereas measurement along the $z$-axis yields opposite spin directions. A pair of electrons described by both (61) and (65) shares the same correlation specified by the spin triplet states in standard quantum mechanics
@@
Similarly, we can give equivalent descriptions of the other two Bell states in standard quantum mechanics. In the case that the initial orientations of angular momentum of electron $A$ are uniformly distributed for all $\theta\in[0,\pi]$, and the correlation between the electron pair are such that $\theta_{A}^{z}=\theta^{z}_{B}$ along the $z$-axis, the joint probability density function is
$\tilde{p}(\theta_{A}^{z},\theta_{B}^{z})=\frac{1}{2}\delta(\theta_{A}^{z})\delta(\theta_{B}^{z})+\frac{1}{2}\delta(\theta_{A}^{z}-\pi)\delta(\theta_{B}^{z}-\pi).$ (67)
@@ V.4 Bell Inequality
Bell inequality is developed to prove that local hidden variable theory cannot produce the predictions of standard quantum mechanics. Bell inequality is violated when tested with an entangled pair of photons or electrons, demonstrating that there is non-local correlation in standard quantum mechanics. Although modern Bell experiments Hensen confirm that such a non-local correlation does not imply non-local causality, this Bell non-locality is still daunting to the quantum physics community because there is no intuitive physical model to explain the phenomenon. Here, we investigate what insights our spin model can bring to this challenge.
@@
Here, $a,a^{\prime}$ are the detector settings for Alice and $b,b^{\prime}$ are the detector settings for Bob. They usually refer to the directions of the magnetic field in the spin measurement and can be considered as unit vectors for the corresponding directions. The term $E(a,b)$ is the expectation value of the measurement outcomes. That is, the statistical average of $S_{A}(a)S_{B}(b)$ where $S_{A}(a)$ and $S_{B}(b)$ are the spin measurement results for setting $a$ and $b$, respectively. For the setting $a$, the unit vector for the direction of the magnetic field is $\hat{a}$, which we still denote as $a$ for simplicity of notation. Then, $S_{A}(a)=+1$ for spin-up and $S_{A}(b)=-1$ for spin-down. Similar meanings can be inferred for the other $E(\cdot)$ terms in (70).
The expectation value $E(\cdot)$ depends on the state of the entangled pair. Suppose that the entangled electron pair is in the singlet described by the joint probability density (61) and (62). It is important to note that both (61) and (62) are needed to give a complete description of the singlet state. In Appendix E, we prove that given the joint probability density (61) and (62),
@@
The reason for the violation of the CHSH inequality is due to the fact that the joint probability density (61) and (62) cannot be factorized. In other words, they cannot be written as a product of two factors, one only contains variable for $A$ and the other only contains variable for $B$. Such a correlation can be maintained even though the two electrons are remotely separated. The root cause of this correlation is that the random variables are correlated, $\theta^{z}_{A}+\theta^{z}_{B}=\pi$ and $\theta^{y}_{A}+\theta^{y}_{B}=\pi$. These random variables are intrinsically local to the electrons. The term non-local correlation, or Bell non-locality, is misleading. Instead, as Hall pointed out Hall2015 , a better term is Bell non-separability.
Figure 7: Bell test with time delay. Starting with a Bell test experiment setup that has already successfully confirmed the violation of Bell-CHSH inequality. Then, we add a time delay in the measurement. Bob delays his measurement by a time period $\Delta t$, while the signals from Alice’s Stern-Gerlach detectors are delayed by the same $\Delta t$ before it is sent to the coincident monitor. When $\Delta t$ is chosen to be sufficiently large, we expect the Bell-CHSH inequality becoming non-violated.
@@ VI Possible Experiments to Confirm the Spin Theory
The spin theory presented here so far recovers many of the results from standard quantum mechanics when certain conditions are taken to the limits. However, when these conditions are not taken to the limits, the present theory will give some results different from those of standard quantum mechanics. In this section, we will provide two possible experiments to confirm the difference.
@@ VI.1 Dependency of Quantization on Interaction with Magnetic Field
In the derivation of (15), one assumes that the parameter $m$, which is related to the order of Tsallis relative entropy as shown in (12), monotonically increases when the electron travels along an inhomogeneous magnetic field $B_{z}$. Standard quantum mechanics postulates that quantization of electron spin is an intrinsic property. Practically, it corresponds to the case that whenever an external field $B_{z}$ is applied, quantization occurs instantaneously, such that only two discrete results can be observed. However, the theory presented here allows $m$ to take a finite number when the magnetic field $B_{z}$ is applied. In that case, the probability density (13) should be used instead of the probability density (15) with two quantized results.
@@
where $^{B}_{0}\hat{z}$ is a constant field, and the coefficient $\eta$ determines the strength of the gradient of the magnetic field. Since the orientation of the intrinsic magnetic moment of the electron follows a probability density distribution (13), the electron passing the Stern-Gerlach apparatus will experience a force that also depends on the orientation. Between angles $\theta$ and $\theta+d\theta$, the force is
@@
The above calculation ignores the effect of the translational component of vacuum fluctuations and only computes the displacement along the $z$-axis using classical mechanics. In theory, this approximation is reasonable when $\Delta T$ is sufficiently large so that $\Delta z$ is much larger than the effect of the translational component of vacuum fluctuations. Eq. (74) shows that the electrons passing through the Stern-Gerlach apparatus will reach the detector screen with a continuous distribution, as shown in Fig. 2b.
However, as $\Delta T$ increases, the parameter $m$ also increases monotonically, the measurement results may rapidly converge to two discrete lines and yield the same prediction as standard quantum mechanics. Such an experiment can be challenging to implement. The next proposed experiment can be more realistic.
@@ VI.2 Bell Test with Time Delay
In Section V.1, we show that for an entangled electron pair described by the joint probability density (61) and (62), the spin correlation can be preserved without a time constraint. However, as explained in Section V.2, when Alice performs her measurement on the spin of electron $A$, obtains spin-up, and makes her prediction on Bob’s measurement result of spin-up along the same direction, the prediction is only valid for a limited time. That is, there is a time constraint specified in (64) that Bob needs to perform his measurement on the spin of the electron $B$ to confirm Alice’s prediction. If $t_{m}^{B}>t_{m}^{A}+\Delta t^{B}_{+}$, Bob will instead obtain the spin-down measurement result. A singlet spin state is described by two probability density functions (61) and (62). After Alice performs her measurement on the electron $A$ and obtains the result of spin-up, if Bob delays his measurement by a time such that the correlation described by either (61) or (62) is no longer satisfied, the singlet state is degraded and no longer valid.
@@
We can explore the effect of entanglement degradation for a singlet state due to Bob’s delayed measurement in the Bell test experiment, as illustrated in Figure 7. In a Bell test experiment, the electron source generates $N$ copies of pairs of entangled electrons in a singlet state. For each pair, electrons $A$ and $B$ are sent to two Stern-Gerlach apparatuses that are remotely separated [^7] As mentioned earlier, in modern Bell test experiment Hensen , the two electrons are already separated remotely before being prepared to be entangled. The electron entanglement is achieved by entanglement swapping with photons. There is no need to move the electrons and, therefore, no need to worry about the impacts on the correlation resulting from the movement of the electrons.. Alice measures the spin of an electron $A$ along an axis labeled by unit vector $a$. At the same time, Bob measures the spin of the electron $B$ along an axis labeled by unit vector $b$. The detectors in each Stern-Gerlach apparatus detect the spin-up and spin-down results and generate the corresponding signals that are sent to the coincidence monitor. Counting the four types (++,+,+,)(++,+-,-+,--) of coincidences for the $N$ copies of the singlet spin pairs, one can calculate whether the Bell-CHSH inequality is violated. As shown in Section V.4, for the singlet spin pair described by (61) or (62), if Alice chooses $a=0$ and $a^{\prime}=\pi/2$, and Bob chooses $b=\pi/4$ and $b^{\prime}=3\pi/4$, the maximum violation of the Bell-CHSH inequality is achieved.
@@
Using exactly the same experiment setup, we add an extra step here. Instead that Bob performs the measurement on electron $B$ at the same time as Alice’s measurement, he delays his measurement by a time period of $\Delta t$. When $\Delta t$ is sufficiently large, the entanglement correlation of a singlet state is degraded, and we expect that the Bell-CHSH inequality will not be violated. Since we do not know how long to delay in order to observe the non-violation of Bell-CHSH inequality, we can start with a very small delay $\Delta t_{1}$ such that the Bell-CHSH inequality is still violated, and then gradually increase the delay. We repeat the experiments $M$ times, but choose delay times such that $\Delta t_{1}<\Delta t_{2}<\ldots<\Delta t_{M}$. Each experiment consumes $N$ copies of singlet electron pairs. According to the present theory, we expect that the violation of Bell-CHSH inequality will be gradually reduced, and at some time the results will satisfy the Bell-CHSH inequality. The reason for this is that with a sufficiently large delay $\Delta t$, the orientation of the intrinsic angular momentum of the electron $B$ can fluctuate to switch from trending upward to trending downward, or vice versa, as depicted in Figure 4. When an orientation switching occurs, either $\theta_{B}\in\Theta_{B}^{+}\to\theta_{B}\in\Theta_{B}^{-}$, or $\theta_{B}\in\Theta_{B}^{-}\to\theta_{B}\in\Theta_{B}^{+}$, the initial correlation between $\theta_{A}$ and $\theta_{B}$ is changed. As $\Delta t$ increases, the orientation of the electron $B$ appears again as a random variable for the $N$ copies of the pairs of electrons, so the initial correlation between $\theta_{A}$ and $\theta_{B}$ no longer holds. Consequently, the Bell-CHSH inequality becomes satisfied. Appendix E gives the mathematical calculation of such a scenario.
@@
The reason why the entanglement relation between $A$ and $B$ is degraded is due to the orientation fluctuations of the intrinsic angular momentum. This is different from the decoherence theory, in which the entanglement relationship between $A$ and $B$ can be destroyed by environmental disturbance. Here, however, the degradation of entanglement during measurement is intrinsic to the electron pair, nothing to do with the environment. Without the measurement event, the entanglement relation between $A$ and $B$ can be preserved without a time limit. Certainly, if there is also environmental disturbance, the entanglement relationship will be destroyed without measurement. So, the present theory is not incompatible with the decoherence theory. However, if confirmed by experiment, the intrinsic degradation of entanglement during measurement can have a non-negligible implication on its applications in quantum computing and quantum information.
@@
The challenges in performing the above experiment come from choosing the appropriate time delay $\Delta t$. To meet the locality condition of the Bell test, $\Delta t$ should be less than the time it takes for a light signal to travel from Alice to Bob. This is to eliminate the potential loophole of hidden signals sent from Alice to Bob. But $\Delta t$ can also not be too small so that the orientation of intrinsic angular momentum of $B$ switches from $\theta_{B}\in\Theta_{B}^{+}\to\theta_{B}\in\Theta_{B}^{-}$, or $\theta_{B}\in\Theta_{B}^{-}\to\theta_{B}\in\Theta_{B}^{+}$. To resolve this contention, one can separate electrons $A$ and $B$ with a very long distance so that the time window to comply with the locality condition is large enough. Denote the time for a light signal to travel between Alice and Bob as $\tau^{AB}$, then the time delay for Bob to perform his measurement must meet the following constraint,
$\tau^{B}_{+}<\Delta t<\tau^{AB}.$ (75)
@@ VII.1 Conceptual Implications
In this paper, an intuitive physical model is provided to understand the phenomenon of entanglement between two spins of electrons. As shown in Section V, entanglement is simply the manifestation of the correlation between the orientations of intrinsic angular momenta of two electrons. Even though the orientations are random variables, they can have a correlation because of previous interactions between them or because of previous interactions with common external fields. Once the correlation of orientations between two electrons is established, it can be maintained even if the two electrons are remotely separated, until a measurement is performed on them. Performing a measurement on one electron will not cause any change on the other electron. However, the measurement outcome reveals the orientation of the measured electron’s intrinsic angular momentum, which enables us to infer the spin measurement outcome of the other electron because of the correlation on the orientations. The seemingly non-local correlation has nothing to do with any non-local causal relation. Although classical correlation can also demonstrate non-local and non-causal relationship, as discussed in Section V.1, the fundamental difference in spin correlation is that the correlated variables (the spin orientations) are intrinsically random.
@@
To further clarify the point that there is no non-causal effect in the Bell experiment with a pair of electrons in the singlet state, suppose after Alice performs her experiment and obtains the result of spin-up for electron $A$, she does not send the measurement outcome to Bob. Since the measurement action of Alice is local and has no influence on the electron $B$, from Bob’s point of view, he still perceives both $A$ and $B$ in the original singlet state described by (61) and (62). Alice’s and Bob’s descriptions on $B$ are different, but both are valid. This is consistent with the relational quantum mechanics (RQM) interpretation advocated by Rovelli Rovelli:1995fv ; Rovelli07 . Since Bob does not know the measurement outcome from Alice, he predicts that if he measures the electron $B$ along any direction, he will obtain spin-up and spin-down with equal probability. However, if Alice sends the measurement results to Bob, Bob can infer the orientation of the intrinsic angular momentum of electron $B$. Bob’s knowledge about electrons $B$ is updated with the new information from Alice. Thus, he can predict that if he measures electron $B$ along the same direction as Alice’s measurement, he will obtain the result of spin-down. This analysis also illustrates that the wave function in quantum mechanics is just a mathematical tool and is not associated with certain ontic properties.
@@
We see that the spin model and theory presented here give an intuitive physical picture that are consistent with several quantum mechanics interpretations. In addition, our theory supports the idea that the measurement outcome is context dependent. The measurement outcome for the spin of an electron depends on the measurement context, that is, the direction of magnetic field used in the Stern-Gerlach apparatus. The two-level discreteness of the measurement outcome is the result of a sufficiently strong gradient of magnetic field. For an electron with initially completely random orientation of intrinsic angular momentum, we can measure the spin along any direction, but always get spin-up and spin-down with equal probability. There is no predefined spin property without specifying an experimental context.